OpenStudy (anonymous):

Would the ? in 16x^2+?x+36 be 6?

OpenStudy (anonymous):

Oops I mean 4.

OpenStudy (anonymous):

36 is the square of ___

OpenStudy (anonymous):

6.

OpenStudy (anonymous):

16 is the square of ___

OpenStudy (anonymous):

4

OpenStudy (anonymous):

4

OpenStudy (anonymous):

\[2\times 6\times 4 =?\]

OpenStudy (anonymous):

48

OpenStudy (anonymous):

you win

OpenStudy (anonymous):

48

OpenStudy (anonymous):

48? But why is it 48?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

because \[(a+b)^2=a^2+2ab+b^2\]

OpenStudy (anonymous):

here \[a=2x\] \[a^2=4x^2\] \[b=6\] \[b^2=36\] and you are looking for \[2ab\]

OpenStudy (anonymous):

if this is a perfect square, than the vales A and C perfectly constrain the values of a and c in Ax^2+Bx+C=(ax+c)^2.

OpenStudy (anonymous):

OH. Thank you. That helps a lot.!

OpenStudy (anonymous):

yw